OpenStudy (anonymous):

HELP WITH DEFERRED PAYMENT PLAN? please Sheila purchased ski equipment for $513 using a six-month deferred payment plan. The interest rate after the introductory period is 18.39%. A down payment of $125 is required as well as a minimum monthly payment of $65. What is the balance at the beginning of the seventh month if only the minimum payment is made during the introductory period?

OpenStudy (amistre64):

do we assume the interest accrues for 6 mnths?

OpenStudy (amistre64):

Month 1 = (513 - 125) * (1+.1839/12) Month 2 = Month 1 * (1+.1839/12) Month 3 = Month 2 * (1+.1839/12) Month 4 = Month 3 * (1+.1839/12) etc ...

OpenStudy (anonymous):

im lost